# Controllability score for linear time-invariant systems on an infinite time horizon

### v1

@@ Appendix B Proof of Lemmas 2 and 3
$y^{\top}\widetilde{W}(p;T)y$ $=\int_{0}^{T}y^{\top}\mathrm{e}^{At}\operatorname{diag}(p)\mathrm{e}^{A^{\top}t}y\mathrm{d}t$
$>0.$
@@
Let $C\in\mathbb{C}^{\ell\times\ell}$ be Hurwitz and $P\in\mathbb{C}^{\ell\times\ell}$ be Hermitian positive definite. Then, the Lyapunov equation
$CY+YC^{*}+P=0$
admits a unique solution $Y\in\mathbb{C}^{\ell\times\ell}$, which is Hermitian positive definite.
@@
We next prove that $W_{0}(p)$ is positive definite. Since $W_{0}(p)$ is a block-diagonal matrix from 11, it suffices to show that $W_{0}^{(\ell)}(p)$ is positive definite for $\ell=1,\ldots,m_{0}^{\prime}$. To simplify the indexing, let $\alpha=m_{-}+m_{0}^{(1)}+\cdots+m_{0}^{(\ell-1)},\ \beta=m_{-}+m_{0}^{(1)}+\cdots+m_{0}^{(\ell)}$. Let $e^{(i)}\coloneq(0,\ldots,0,1)^{\top}\in\mathbb{R}^{n_{i}\times n_{i}}$. Then, a direct calculation yields
@@
$E^{(\ell)}$ $\coloneq\begin{pmatrix}e^{(\alpha+1)}&0&\cdots&0&0\\ 0&e^{(\alpha+2)}&\cdots&0&0\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&\cdots&e^{(\beta-1)}&0\\ 0&0&\cdots&0&e^{(\beta)}\end{pmatrix}$
$\in\mathbb{R}^{n_{0}^{(\ell)}\times m_{0}^{(\ell)}},$
@@ Appendix C Proof of Theorems 2 and 3
Let $\{f^{(k)}\}_{k\in\mathbb{N}}$ be a sequence of functions $f^{(k)}\colon\mathbb{R}^{n-1}\to\mathbb{R}\cup\{\infty\}$. We say that $\{f^{(k)}\}_{k\in\mathbb{N}}$ is equi level-bounded if there exists a bounded set $B\subset\mathbb{R}^{n-1}$ such that $\mathrm{lev}_{\leq\alpha}f^{(k)}\subset B$ for all $k$. We say that $\{f^{(k)}\}_{k\in\mathbb{N}}$ epi-converges to $f$ if, for every $p^{\prime}\in\mathbb{R}^{n-1}$,
@@
Proposition 3 ([21, Theorem 7.33])
Assume that $\{f^{(k)}\}_{k\in\mathbb{N}}$ is an equi level-bounded sequence of functions $f^{(k)}\colon\mathbb{R}^{n-1}\to\mathbb{R}\cup\{\infty\}$, that $\{f^{(k)}\}_{k\in\mathbb{N}}$ epi-converges to $f$, and that $f^{(k)}$ and $f$ are lower semicontinuous and proper. Then, the sequence $\{\widetilde{p}^{(k)}\}_{k\in\mathbb{N}}$ such that $p^{(k)}\in\operatorname*{arg\,min}f^{(k)}$ is bounded, and all its cluster points lie in $\operatorname*{arg\,min}f$. In particular, when $\operatorname*{arg\,min}f=\{\widetilde{p}^{*}\}$, the sequence $\{\widetilde{p}^{(k)}\}_{k\in\mathbb{N}}$ converges to $\widetilde{p}^{*}$.
@@
Let $C_{1},C_{2}\in\mathbb{C}^{\ell\times\ell}$ be Hermitian positive definite matrices, and $\mu>0$ is the minimum eigenvalue of $C_{2}$. Then,
$\operatorname{tr}(C_{1}C_{2})\geq\mu\operatorname{tr}(C_{1})$
holds.
@@
Since $C_{2}\succeq\mu I_{\ell}$, we obtain $C_{1}^{1/2}C_{2}C_{1}^{1/2}\succeq\mu C_{1}$. Therefore,
$\operatorname{tr}(C_{1}C_{2})=\operatorname{tr}(C_{1}^{1/2}C_{2}C_{1}^{1/2})\geq\mu\operatorname{tr}(C_{1})$
holds. $\Box$
@@
and therefore $\widetilde{p}\in\mathrm{lev}_{\leq\alpha}h$. This shows that $\mathrm{lev}_{\leq\alpha}h$ is closed, implying that $h$ is lower semicontinuous. Similarly, we can prove that $h^{(k)}$ is lower semicontinuous. The proof proceeds in the same manner as that of Theorem 2. $\Box$
@@ Appendix D Proof of Theorems 4 and 5
To prove Theorems 4 and 5, we prepare the following lemmas, which can be proved in the same manner as [23, Theorems 1 and 3], although Problems 2 and 6 have slightly different forms.
@@
where we set $\alpha=m_{-}+m_{0}^{(1)}+\cdots+m_{0}^{(\ell-1)},\ \beta=m_{-}+m_{0}^{(1)}+\cdots+m_{0}^{(\ell)}$ to simplify the indexing.
@@
Let $x=(x_{1},\ldots,x_{n})^{\top}\in\mathbb{R}^{n}$ be any vector such that $W_{-}(x)=0$. Then, from Lemma 8, it suffices to prove $x=0$. Let $\widetilde{X}\coloneq Q^{-1}XQ^{-*}$ and partition $\widetilde{X}$ in the same way as 30. By an argument analogous to that of Theorem 4, we obtain $\widetilde{X}_{-,-}=O$. Proceeding as in the proof of Theorem 4, the assumption then implies that $x=0$, which completes the proof. $\Box$
@@ Appendix E Uniqueness of finite-horizon controllability scores
Moreover, we employ the following lemma, which can be proved in the same way as in [23, Theorems 1 and 3].
@@
It follows from 31 that
$\operatorname{vec}\left(\widetilde{W}(B;T)\right)$ $=\int_{0}^{T}\operatorname{vec}\left(\mathrm{e}^{At}B\mathrm{e}^{A^{\top}t}\right)\mathrm{d}t$
$=\left(\int_{0}^{T}\mathrm{e}^{At}\otimes\mathrm{e}^{At}\mathrm{d}t\right)\operatorname{vec}(B)$
$=\left(\int_{0}^{T}\mathrm{e}^{Kt}\mathrm{d}t\right)\operatorname{vec}(B).$
@@
which is a holomorphic function. By identifying 32 and 33, we have $F_{T}=f_{T}(K)$. Consequently, from Lemmas 9 and 10, every eigenvalue of $F_{T}$ is of the form $f_{T}(\lambda_{i}+\lambda_{j})$ for some $i,j$.

### v2

@@ Appendix B Proof of Lemmas 2 and 3
$y^{\top}\widetilde{W}(p;T)y=\int_{0}^{T}y^{\top}\mathrm{e}^{At}\operatorname{diag}(p)\mathrm{e}^{A^{\top}t}y\mathrm{d}t>0.$
@@
Let $C\in\mathbb{C}^{\ell\times\ell}$ be Hurwitz and $P\in\mathbb{C}^{\ell\times\ell}$ be Hermitian positive definite. Then, the Lyapunov equation $CY+YC^{*}+P=O$ admits a unique solution $Y\in\mathbb{C}^{\ell\times\ell}$, which is Hermitian positive definite.
@@
We next prove that $W_{0}(p)$ is positive definite. Since $W_{0}(p)$ is a block-diagonal matrix from 11, it suffices to show that $W_{0}^{(\ell)}(p)$ is positive definite for $\ell=1,\ldots,m_{0}^{\prime}$. To simplify the indexing, let $\alpha=m_{-}+m_{0}^{(1)}+\cdots+m_{0}^{(\ell-1)},\ \beta=\alpha+m_{0}^{(\ell)}$. Let $e^{(i)}\coloneq(0,\ldots,0,1)^{\top}\in\mathbb{R}^{n_{i}\times n_{i}}$. Then, a direct calculation yields
@@
$E^{(\ell)}\coloneq\begin{pmatrix}e^{(\alpha+1)}&\cdots&0\\ \vdots&\ddots&\vdots\\ 0&\cdots&e^{(\beta)}\end{pmatrix}\in\mathbb{R}^{n_{0}^{(\ell)}\times m_{0}^{(\ell)}},$
@@ Appendix C Proof of Theorems 2 and 3
Before proving Theorems 2 and 3, we collect several definitions and preliminary facts that will be used throughout the proofs.
Let $\widetilde{\Delta}$ be the simplex in $\mathbb{R}^{n-1}$, i.e.,
The mapping $\varphi\colon\widetilde{\Delta}\to\Delta$ defined by
is a homeomorphism; its inverse is given by
$\varphi^{-1}(p_{1},\ldots,p_{n})=(p_{1},\ldots,p_{n-1})^{\top}.$
For a set $\Omega\subset\mathbb{R}^{n-1}$, $\operatorname*{int}\Omega$ denotes its interior.
@@
Let $\{f^{(k)}\}_{k\in\mathbb{N}}$ be a sequence of functions $f^{(k)}\colon\mathbb{R}^{n-1}\to\mathbb{R}\cup\{\infty\}$. We say that $\{f^{(k)}\}_{k\in\mathbb{N}}$ is equi level-bounded if for every $\alpha$, there exists a bounded set $B\subset\mathbb{R}^{n-1}$ such that $\mathrm{lev}_{\leq\alpha}f^{(k)}\subset B$ for all $k$. We say that $\{f^{(k)}\}_{k\in\mathbb{N}}$ epi-converges to $f$ if, for every $\widetilde{p}\in\mathbb{R}^{n-1}$,
@@
Proposition 3 ([22, Theorem 7.33])
Assume that $\{f^{(k)}\}_{k\in\mathbb{N}}$ is an equi level-bounded sequence of functions $f^{(k)}\colon\mathbb{R}^{n-1}\to\mathbb{R}\cup\{\infty\}$, that $\{f^{(k)}\}_{k\in\mathbb{N}}$ epi-converges to $f$, and that $f^{(k)}$ and $f$ are lower semicontinuous and proper. Then, the sequence $\{\widetilde{p}^{(k)}\}_{k\in\mathbb{N}}$ such that $\widetilde{p}^{(k)}\in\operatorname*{arg\,min}f^{(k)}$ is bounded, and all its cluster points lie in $\operatorname*{arg\,min}f$. In particular, when $\operatorname*{arg\,min}f=\{\widetilde{p}^{*}\}$, the sequence $\{\widetilde{p}^{(k)}\}_{k\in\mathbb{N}}$ converges to $\widetilde{p}^{*}$.
@@
Let $C_{1},C_{2}\in\mathbb{C}^{\ell\times\ell}$ be Hermitian positive definite matrices, and $\mu>0$ is the minimum eigenvalue of $C_{2}$. Then, $\operatorname{tr}(C_{1}C_{2})\geq\mu\operatorname{tr}(C_{1})$ holds.
@@
Since $C_{2}\succeq\mu I_{\ell}$, we obtain $C_{1}^{1/2}C_{2}C_{1}^{1/2}\succeq\mu C_{1}$. Therefore, $\operatorname{tr}(C_{1}C_{2})=\operatorname{tr}(C_{1}^{1/2}C_{2}C_{1}^{1/2})\geq\mu\operatorname{tr}(C_{1})$. $\Box$
@@
and therefore $\widetilde{p}\in\mathrm{lev}_{\leq\alpha}h$. This shows that $\mathrm{lev}_{\leq\alpha}h$ is closed, implying that $h$ is lower semicontinuous. Similarly, we can prove that $h^{(k)}$ is lower semicontinuous.
As $T\to\infty$, we have $D(T)^{-1}\to\begin{pmatrix}I_{n_{-}}\\ &O\end{pmatrix}$, and Theorem 1 implies that $D(T_{k})^{-*}W(\varphi(\widetilde{p});T_{k})^{-1}D(T_{k})^{-1}\to\begin{pmatrix}W_{-}(\varphi(\widetilde{p}))^{-1}\\ &O\end{pmatrix}$. Therefore, the remainder of the proof follows the same steps as in the proof of Theorem 2. $\Box$
@@ Appendix D Proof of Theorems 4 and 5
To prove Theorems 4 and 5, we prepare the following lemmas, which can be proved in the same manner as [24, Theorems 1 and 3], although Problems 2 and 6 have slightly different forms.
@@
where we set $\alpha=m_{-}+m_{0}^{(1)}+\cdots+m_{0}^{(\ell-1)},\ \beta=\alpha+m_{0}^{(\ell)}$ to simplify the indexing.
@@
Let $x=(x_{1},\ldots,x_{n})^{\top}\in\mathbb{R}^{n}$ be any vector such that $W_{-}(x)=0$. Then, from Lemma 8, it suffices to prove $x=0$. Let $\widetilde{X}\coloneq Q^{-1}XQ^{-*}$ and partition $\widetilde{X}$ in the same way as 31. By an argument analogous to that of Theorem 4, we obtain $\widetilde{X}_{-,-}=O$. Proceeding as in the proof of Theorem 4, the assumption then implies that $x=0$, which completes the proof. $\Box$
@@ Appendix E Uniqueness of finite-horizon controllability scores
Moreover, we employ the following lemma, which can be proved in the same way as in [24, Theorems 1 and 3].
@@
It follows from 32 that
$\operatorname{vec}\left(\widetilde{W}(B;T)\right)=\int_{0}^{T}\operatorname{vec}\left(\mathrm{e}^{At}B\mathrm{e}^{A^{\top}t}\right)\mathrm{d}t\\ =\left(\int_{0}^{T}\mathrm{e}^{At}\otimes\mathrm{e}^{At}\mathrm{d}t\right)\operatorname{vec}(B)=\left(\int_{0}^{T}\mathrm{e}^{Kt}\mathrm{d}t\right)\operatorname{vec}(B).$
@@
which is a holomorphic function. By identifying 33 and 34, we have $F_{T}=f_{T}(K)$. Consequently, from Lemmas 9 and 10, every eigenvalue of $F_{T}$ is of the form $f_{T}(\lambda_{i}+\lambda_{j})$ for some $i,j$.